Two cyclists leave a city at the same time, one going east the other going west. The west bound cyclist bikes 4 mph faster than the east bound cyclist. After 5 hours they are 200 miles apart. How fast is the east bound cyclist riding?
Found 2 solutions by solver91311, Finavon:
Answer by solver91311(24713)



The sum of their speeds had to be 200 divided by 5, so:



Solve for

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by Finavon(81)

speeds e and w
w=e+4
in one hour they will be 200/5 miles apart
w+e=40
substitute for w
e+4+e=40
2e=36
e=18
w=18+4=22
East at 18 mph; West at 22 mph